OpenStudy (anonymous):

Which of the following is a factor of 2x^4 + 22x^3 + 60x^2? 2x^3 x^4 x + 4 x + 5

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

GCF... 2 and x^2

OpenStudy (anonymous):

sooo... 2x^4 + 22x^3 + 60x^2 2 and x^2 2x^2(x^2 + 11x^1 + 30)

ganeshie8 (ganeshie8):

perfect!

OpenStudy (anonymous):

lol weird formula time xD

ganeshie8 (ganeshie8):

next we need to factor x^2 + 11x + 30

ganeshie8 (ganeshie8):

formulas wont work this time :o

OpenStudy (anonymous):

o_O

OpenStudy (anonymous):

lol what do we do then

ganeshie8 (ganeshie8):

this is a new method, see if u can follow it :- first, we pick two numbers \(a\), \(b\) such that :- \(a*b\) = 30 \(a+b\) = 11

ganeshie8 (ganeshie8):

can u tell me the two numbers ?

OpenStudy (anonymous):

10 * 3 = 30 5 + 6 = 11

ganeshie8 (ganeshie8):

noo, both shud be same

ganeshie8 (ganeshie8):

a*b = 30 and, same a+b = 11

ganeshie8 (ganeshie8):

like that...

ganeshie8 (ganeshie8):

5*6 = 30 5+6 = 11

ganeshie8 (ganeshie8):

so, 5 and 6 are our numbers, ok

OpenStudy (anonymous):

alright I got it

ganeshie8 (ganeshie8):

next :- 2x^2(x^2 + \(11x\) + 30) 2x^2(x^2 + \(5x + 6x\) + 30)

ganeshie8 (ganeshie8):

fine ? i just wrote 11x as 5x+6x

OpenStudy (anonymous):

yeah im following ya

ganeshie8 (ganeshie8):

2x^2(x^2 + \(11x\) + 30) 2x^2(x^2 + \(5x+6x\) + 30) -------- --------

ganeshie8 (ganeshie8):

factor those two separately

ganeshie8 (ganeshie8):

2x^2(x^2 + \(11x\) + 30) 2x^2(x^2 + \(5x+6x\) + 30) -------- -------- 2x^2(x(x+\(5\))+\(6\)(x + 5)) 2x^2(x+\(5\))(x + \(6\))

ganeshie8 (ganeshie8):

take it slow... if it doesnt make fully sense its okay, you need to practice 2-3 problems to see how the method works

OpenStudy (anonymous):

Yeah i understand it a bit, thanks :)

ganeshie8 (ganeshie8):

good to hear :)